### v1

# On the Roles of LLMs in Planning: Embedding LLMs into Planning Graphs

Abstract
Plan synthesis aims to generate a course of actions or policies to transit given initial states to goal states, provided domain models that could be designed by experts or learnt from training data or interactions with the world. Intrigued by the claims of emergent planning capabilities in large language models (LLMs), works have been proposed to investigate the planning effectiveness of LLMs, without considering any utilization of off-the-shelf planning techniques in LLMs. In this paper, we aim to further study the insight of the planning capability of LLMs by investigating the roles of LLMs in off-the-shelf planning frameworks. To do this, we investigate the effectiveness of embedding LLMs into one of the well-known planning frameworks, graph-based planning, proposing a novel LLMs-based planning framework with LLMs embedded in two levels of planning graphs, i.e., mutual constraints generation level and constraints solving level. We empirically exhibit the effectiveness of our proposed framework in various planning domains.
1 Introduction
Plan synthesis aims to generate a course of actions or policies to transit given initial states to goal states, provided domain models that could be designed by experts or learnt from training data (Aineto et al., 2019) or interactions with the world (Lamanna et al., 2023; Jin et al., 2022b). It is a time- and space-consuming open issue in the planning community (Ghallab et al., 2004). Intrigued by the claims of emergent planning capabilities in large language models (LLMs), works have been proposed to investigate the planning effectiveness of LLMs, without considering any utilization of off-the-shelf planning techniques in LLMs (Valmeekam et al., 2023b). As demonstrated by (Valmeekam et al., 2023b), even in a seemingly simple common-sense domain like Blocksworld that humans usually find easy to solve, LLMs are evaluated to be quite ineffective in planning autonomously.
An interesting result shown by (Valmeekam et al., 2023b) is when taking the solution generated by LLMs, which is generally incorrect, as a seed plan to be repaired by an off-the-shelf planner, e.g., LPG (Gerevini & Serina, 2002), a significant improvement in search steps can be seen over the empty plan provided as a seed plan for the planner. This indicates that LLMs can indeed provide some helpful information (e.g., in some sense of heuristics) for planning, even though they cannot solve planning problems solely. Inspired by the result of loosely using plans generated by LLMs as seed plans, we are curious if it is possible to “dig” more helpful information from LLMs to assist planning deeply, e.g., by inserting LLMs into planning frameworks. By doing this, we aim to answer the question: what roles can be played exactly by LLMs in planning? Indeed, there have been attempts to explore off-the-shelf planning techniques to help LLMs solving planning problems (Liu et al., 2023). Similar to (Valmeekam et al., 2023b), they only view planners as black-boxes without deepening the integration of LLMs in planning frameworks.
To do this, we investigate the effectiveness of embedding LLMs into one of the well-known planning frameworks, graph-based planning (Blum & Furst, 1997). We propose a novel LLMs-based planning framework with LLMs embedded in two levels of the planning framework (namely LLMs4Plan), i.e., (a) selecting promising actions in each action-level to expand the planning graph, and (b) building non-mutual action sets in each action-level of the expanded planning graph, which correspond to two critical steps that influence the efficiency and effectiveness of graph planning.
(a) Expantion
(b) Backtracking
Figure 1: Two critical steps in graph planning
For example, as shown in Figure 1(a), there could be many actions in “action-level 1” after expanding the planning graph based on “state-label 0” with Graphplan (Blum & Furst, 1997). We expect that LLMs will help select only a few promising actions in each action level, e.g., $\{a_{1},a_{2},a_{3}\}$ in “action-level 1”. In Figure 1(b), when backtracking from “state-level $K$” that contains goals, there could be many candidate sets of actions to be explored (e.g., “action set 1”, “action set 2”, “action set 3”) — actions in each candidate set are not mutual with each other. It is particularly time-consuming to search all of the valid candidate sets based on the mutual constraints in each action-level. We expect that LLMs are capable of helping select only a few candidate sets to be backtracked, e.g., only “action set 1” is selected to be backtracked by LLMs.
Specifically, in LLMs4Plan, for each action-level $i$, we first automatically generate prompts based on propositions in state-level $i-1$, goals and domain models when using LLMs to select actions to expand planning graphs in action-level $i$. After that, when backtracking from the last state-level of expanded planning graphs, i.e., the last state-level in expanded planning graphs contains goals, we automatically generate prompts based on propositions in state-level $j$ (from which backtracking is conducted), propositions from state-level 0, mutual constraints in action-level $j$, and domain models. We embed the above two components into one of the well-known off-the-shelf graph planners, Graphplan. We study the effectiveness of different cases of adding or removing the above one or two components in Graphplan to see the significance of roles LLMs play in the graph planning framework.
Through this study, we provide new clues for how to deeply embed LLMs into off-the-shelf planning frameworks, i.e., first identifying critical steps (generally time-consuming ones) in specific planning frameworks, and then designing proper prompt generation procedure to be embedded into the frameworks. We verify that soly relying on LLMs to do planning is far from a good option, while leveraging LLMs to help deal with some critical steps in the graph planning framework is possible.
2 Related work
LLMs as planners: Many works have investigated how to leverage LLMs to accomplish planning tasks, such as Chain-of-Thoughts (Wei et al., 2022), Tree-of-Thoughts (Yao et al., 2023a) and Zero-Shot Planner (Huang et al., 2022a), which utilize prompts to guide LLMs for generating action sequences for planning related tasks. Approaches such as HuggingGPT (Shen et al., 2023) and Chameleon (Lu et al., 2023) aim to generate initial plans using different tools and then call the corresponding APIs for execution by augmenting LLMs with plug-and-play modules for compositional reasoning. They designed an LLM-based planner to assemble a sequence of tools to execute to generate final solution plans for planning tasks. There are also some works that prompt LLMs to compose plans in the form of PDDL (Planning Domain Definition Language) in planning community (Liu et al., 2023; Pallagani et al., 2023). All of the above-mentioned approaches generally synthesize plans without consideration of feedback from external environments. In order to consider environmental feedback, SayCan (Ichter et al., 2022), ReAct (Yao et al., 2023b), Reflexion (Shinn et al., 2023) and Inner Monologue (Huang et al., 2022b) allow LLMs to take single-step actions according to the environmental feedback, assuming LLM-generated initial plans are correct without any adaptation of them. They only adjust the immediate action being executed and are prone to fall into local sub-optimal actions without considering long-term goals. In consideration of removing this assumption, the evaluation of planning capabilities in the literature involves the user incrementally interacting with LLMs, and re-prompting it to point out flaws in its plans, with the hope that the LLM eventually reaches an executable plan (Huang et al., 2022b; Yao et al., 2023b; Raman et al., 2022; Wang et al., 2023; Sun et al., 2023; Silver et al., 2023). Such evaluations are notoriously with the actual planning being done by humans in the loop rather than LLMs themselves. Instead of considering human-in-the-loop, (Valmeekam et al., 2023a) built another framework for automatically evaluating planning capabilities of LLMs by leveraging automated planning models and tools to generate the queries and validate answers from LLMs. All of the above-mentioned approaches view LLMs as a sole planner rather than being leveraged as components and embedded into off-the-shelf planning frameworks, which means the planning capability of planning frameworks is not leveraged by them.
LLMs finetuned to be Planners: Another issue of broad interest in LLMs is to investigate how to finetune LLMs to be capable of planning. Some works (Reed et al., 2022; Brohan et al., 2023) collected paired visual, language and action data and trained an enormous neural networks for tackling long-horizon tasks. Collecting those data, however, is often expensive and difficult to scale up. Different from training LLMs from scratch, finetuning off-the-shelf LLMs (Driess et al., 2023; Li et al., 2022) to deal with visual and language related tasks using task-specific robot demonstrations is an option to alleviate issues of data collection. It is, however, still problematic since finetuning most high-performing language models, such as GPT3.5/4 (Ouyang et al., 2022) and PaLM (Chowdhery et al., 2023), is currently impossible, as the model weights are not open-sourced. Besides, paired vision and language robotics demonstrations, which are still needed for finetuning, are not readily available in practice. Considering this, (Ajay et al., 2023) propose compositional foundation models for hierarchical planning, where a foundation model is a composition of different expert models trained on language, vision, and action data individually. Given an abstract language instruction describing the desired task, they use a large language model to generate a sequence of sub-tasks and then use a large video diffusion model to capture geometric and physical information about the world and generate a more detailed plan. Finetuning LLMs to be a planner indicates training data are required to be available and the planning capability of finetuned LLMs to be transferred to other tasks is still an open issue. Instead of finetuning LLMs, we aim to leverage both benefits of LLMs and planning frameworks to solve planning tasks without any training data.
LLMs as components in planners: Some studies try to use LLMs as heuristics (Zhang et al., 2023) or transition function (Hao et al., 2023) in MCTS, boosting the performance in coding or small-scale reasoning. Utilizing LLMs as a world model in depth, however, is not discussed (Zhao et al., 2023). LLMs can also provide a commonsense model of the world in addition to a policy that acts on it. The world model and the policy are combined in Monte Carlo Tree Search (MCTS) to scale up task planning (Zhao et al., 2023), which is demonstrated to be effective by embedding LLMs in the MCTS framework. The idea is similar to our work in the sense that LLMs are embedded into an off-the-shelf framework in depth. However, they aim to learn MCTS policies from interactions with environments, while our work aims to solve planning problems with LLMs embedded into an off-the-shelf planner, planning-graph planner, without any interactions or learning from environments.
3 Problem Formulation
In this work we consider classical planning problems specified in the form of STRIPS language (Fikes & Nilsson, 1971). Similar idea can be extended into more expressive planning language such as PDDL (Ghallab et al., 1998). A STRIPS domain is composed of a set of action models $\mathcal{A}$, each of which is a quadruple $\langle a,\textrm{PRE}(a),\textrm{ADD}(a),\textrm{DEL}(a)\rangle$, where $a$ is an action name with zero or more parameters, $\textrm{PRE}(a)\subseteq\mathcal{L}$ is a precondition list indicating the conditions under which $a$ can be applied, $\textrm{ADD}(a)\subseteq\mathcal{L}$ is an adding list and $\textrm{DEL}(a)\subseteq\mathcal{L}$ is a deleting list indicating the effects of $a$. $\mathcal{L}$ is a set of atoms, each of which is composed of a predicate with zero or more parameters. We define a planning problem as $\mathcal{P}=\langle\mathcal{R},s_{0},g,\mathcal{A}\rangle$, where $s_{0}\subseteq\mathcal{R}$ is an initial state and $g\subseteq\mathcal{R}$ is a goal. $\mathcal{R}$ is a set of propositions. A solution plan $\pi=\langle A,C\rangle$ to $\mathcal{P}$ consists of a set of actions $A$ and a set of partial order constraints $C$ among actions. Some actions in $A$ are allowed to be executed in parallel so long as there are no conflicts among them, e.g., there are conflicts between two actions if one deletes a precondition or an adding effect of the other (more detailed description of conflicts will be given in Section 4.2). A partial order plan can be equivalently stretched into a total order linear plan, which means both of them can successfully transit initial state $s_{0}$ to goal $g$. An intuitive example of our planning problem is as shown below.
=========== Beginning of the example ============
Example 1: Suppose we would like to clean a bedroom using a vacuum which is placed in a tool room. We can formulate the problem $\mathcal{P}=\langle\mathcal{R},s_{0},g,\mathcal{A}\rangle$ in the form of planning language as shown below (note that we assume there is no parameter for each predicate and action for simplicity):
  • the set of propositions: =$\mathcal{R}=\{dirty(),toolroom(),clean(),bedroom()\}$, which indicates the set of all propositions in the problem.
  • initial state: $s_{0}$={dirty(), toolroom()}, which indicates the bedroom to be cleaned is dirty, and the tool vacuum is in the tool room.
  • goal: $g=\{clean(),toolroom()\}$, which indicates the bedroom is clean, and the tool vacuum is back to the tool room.
  • action models: $\mathcal{A}$ is as shown below:
    Action Preconditions Effects
    $vacuum()$ $dirty()$, $clean()$,
    $bedroom()$$\neg dirty()$
    $move2tr()$ $bedroom()$ $toolroom()$,
    $\neg bedroom()$
    $move2br()$ $toolroom()$ $bedroom()$,
    $\neg toolroom()$
    Action vacuum aims to vacuuming the bedroom, the preconditions of which are the bedroom is dirty and the vacuum-cleaner is in the bedroom. The effects of vacuum are the bedroom is clean (deleting dirty() as well). Action $move2tr$ aims to move the vacuum-cleaner to the tool room, the precondition of which is the vacuum-cleaner is in the $bedroom$. The effects are the vacuum-cleaner is in the tool room and not in the bedroom. Similarly, action $move2br$ aims to move the vacuum-cleaner to the bedroom, the precondition of which is the vacuum-cleaner is in the tool room. The effects are the vacuum-cleaner is in the bedroom and not in the tool room.
The solution plan $\pi$ to the problem $\mathcal{P}$ is: $move2br(),vacuum(),move2tr()$.
============= End of the example ==============
4 Our LLMs4Plan approach
An overview of our LLMs4Plan approach is as shown in Algorithm 1. In Step 1, we first build a planning graph $PG$ with the help of LLMs given the planning problem $\mathcal{P}$ such that the goal is included by the last state-level in $PG$. If the number of levels in $PG$ is less than the maximal levels $MaxIter$, we build mutual constraints $\mathcal{C}$ and sort the candidate sets of actions, resulting in the list of candidate sets $\mathbb{A}$ — each of which satisfies $\mathcal{C}$ — with the help of LLMs. After that we make a deep first search based on $\mathbb{A}$ to get $\pi$. If $\pi$ is not Failure, it is returned as the solution plan; otherwise, we expand $PG$ with one more level using LLMs and repeat Steps 2-11 of the algorithm until the maximal level $MaxIter$ is reached or a solution plan $\pi$ is found and returned in Step 7. In the following subsections, we will detail main steps of Algorithm 1.
Algorithm 1 An overview of our LLMs4Plan approach
Input: Planning problem $\mathcal{P}$
Output: Solution $\pi$
1:  Build graph: $(PG,k)=buildGraphLLMs(\mathcal{P})$
2:  while $k<MaxIter$ do
3:     Build constraints: $\mathcal{C}=buildConstraints(PG)$
4:     Sort action sets: $\mathbb{A}=sortActionsLLMs(PG,\mathcal{C})$
5:     Search solutions: $\pi=depthFirstSearch(\mathbb{A},PG)$
6:     if $\pi\neq$ Failure then
7:        return $\pi$
8:     end if
9:     Expand graph: $PG\leftarrow expandGraphLLMs(PG)$
10:     $k\leftarrow k+1$
11:  end while
12:  return Failure
4.1 Building Planning Graphs with LLMs
A planning graph $PG$ is the search space for a relaxed version of the planning problem, an intuitive framework of which is shown in Figure 5. It alternates layers of ground literals and actions. “Square” nodes at action-level $i+1$ indicate actions that might be possible to be executed in state $s_{i}$. Maintenance actions indicate dump operators that keep literals unchanged between state-levels $i$ and $i+1$. “Black circle” nodes at state-level $i$ indicate literals that might possibly be true at time $i$. Edges between state-level $i$ and action-level $i$ indicate literals in state-level $i$ are preconditions of actions in action-level $i$, while edges between action-level $i$ and state-level $i+1$ indicate literals in state-level $i+1$ are adding or deleting effects of actions in action-level $i$. The nodes in the first state-level indicate literals that are true in initial state $s_{0}$.
Figure 2: The framework of a planning graph
The procedure of build the planning graph with LLMs (i.e., buildGraphLLMs) based on the given planning problem $\mathcal{P}=\langle\mathcal{R},s_{0},g,\mathcal{A}\rangle$ is as follows:
  1. 1.
    All propositions in $s_{0}$ and the negation of propositions in $\mathcal{R}-s_{0}$ are added into state-level 0.
  2. 2.
    All actions in $\mathcal{A}$, whose preconditions are satisfied in state-level 0 and selected by LLMs, are added into action-level 1; a maintenance action corresponding to each proposition in state-level 0 is added into action-level 1.
  3. 3.
    The propositions added or deleted by actions in action-level 1 are added into state-level 1; and all propositions in state-level 0 are added into state-level 1 as well (i.e., which is done by the maintenance action).
  4. 4.
    We repeat steps 1-3 by increasing state-level 0 to 1 (or $i$ to $i+1$) until all propositions in goal $g$ are included by state-level $k$.
Figure 3: The prompt for consulting LLMs
[…]
4.2 Building Mutual Constraints
Due to the satisfaction of action models being relaxed, actions and/or states in action-levels or state-labels may be inconsistent, i.e., there may be some actions mutually exclusive in action-levels, or some literals mutually exclusive in state-levels. As shown in Figure 4, there are there types mutual exclusion constraints among actions.
Figure 4: Mutual exclusion of actions
Specifically, two actions at the same action-level are mutex, if they satisfy the following conditions:
  • An effect of one negates an effect of the other, which is called inconsistent effects.
  • One deletes a precondition of the other, which is called interference.
  • They have mutually exclusive preconditions, which is called Competing needs.
Otherwise they do not interfere with each other, i.e., both may appear in a solution plan. Two literals at the same state-level are mutex if one is the negation of the other, or all ways of achieving them are pairwise mutex, namely inconsistent support.
An example planning graph corresponding to Example 1 is as shown in Figure 5.
Figure 5: An example of planning graph and mutual constraints indicated in RED arcs
Action vacuum is mutually exclusive with action dumb for toolroom at action-level 2 since vacuum’s precondition bedroom is mutually exclusive with toolroom at state-level 1.
4.3 Sort Action Sets with LLMs and Search Solutions
After we build the set of constraints in Step 3 of Algorithm 1, we use off-the-shelf procedure presented in (Blum & Furst, 1997) to compute candidate action sets such that there are no conflicts (i.e., satisfying the constraints $\mathcal{C}$) among actions in each action set. After that, we consult LLMs to sort the action sets by designing the prompts as shown in Figure 6, which is similar to the prompt shown in Figure 3 except the command in BLUE. After we get the sorted action sets $\mathbb{A}$, we conduct the dept-first search procedure as done in (Blum & Furst, 1997) by giving the priority of action sets based on the sorted action sets in $\mathbb{A}$.
Figure 6: The prompt for consulting LLMs to sort action sets
In Step 9 of Algorithm 1 (i.e., $expandGraphLLMs(PG)$), we expand $PG$ with the same prompt as shown in Figure 3 when selecting actions for expanding one more level of $PG$.
5 Experiment
5.1 Experimental Setup
In this experiment, we pick four domains with different scenarios, including gripper, miconic, logistics, and movie. 10 problems are randomly selected for each domain. The specific scenarios and scale are described as follows:
  • Gripper: Tasks utilize an existing robot arm to move objects between rooms, with between 6 and 20 objects in the scene.
  • Miconic: Tasks uses elevators to serve guests between floors and help them reach the floor they want to go to, with between 6 and 50 objects in the scene.
  • Logistics: One of the classic logistics problems. Transportation of items between different locations in different cities using trucks and airplanes, with between 10 and 30 objects in the scenario.
  • Movie: Simulate some simple behaviors while watching a movie with between 45 and 155 objects in the scene.
To demonstrate the effectiveness of our LLMs4Plan approach, we designed five sets of comparison experiments. The methods were implemented using Python. We compared our LLMs4Plan approach with four other methods, which are listed below:
  • GP: It is the graph-based planning algorithm mentioned above. We implement the traditional graph planning algorithm as the most important baseline for comparison. We directly provide domain.pddl and problem.pddl to the planner for solving.
  • GPT-3.5: We simply construct and splice the contents of domain.pddl and problem.pddl directly. We then add the necessary command prompts to form a complete prompt into GPT3.5 and command the model to solve the problem directly.
  • GPT-4: The process is the same as GPT-3.5.
  • LLMs4Plan-GPT3.5: When we are expanding the hierarchy or backtracking, we are faced with a candidate action selection with LLMs. We guide the LLMs to select a minimal subset of actions from them and utilize this subset of actions for the next algorithmic operations, where LLMs are specifically GPT-3.5.
  • LLMs4Plan-GPT4: Replaced the LLM model with GPT4, otherwise same as LLMs4Plan-GPT3.5.
5.2 Experimental Results
In the experimental framework described, we employed five distinct metrics to assess the efficacy of various methodologies: the problem-solving success rate, the cumulative count of expansion actions, the aggregate number of mutually exclusive actions, the quantity of expansion layers, and the node count for backtracking in Depth-First Search (DFS).
Problem-solving success rate. The solvability of a problem is a crucial metric in assessing planning problems. All approaches are required to generate a sequence of actions that is sufficient to solve the problem, and only if the problem can be transferred from the initial state to the goal state through this sequence of actions can the corresponding problem be considered to be successfully solved. Furthermore, we imposed a maximum time constraint of five minutes for problem resolution; any problems unsolved within this duration are deemed failures.
Total number of expansion actions. In the GP algorithm, the expansion of actions at each layer is a fundamental process, and the number of these expansions serves as a vital metric. Under the premise of preserving effective actions, fewer expansions result in a reduced count of mutually exclusive action pairs and subsequently fewer branches in the deep search phase of backtracking, thereby enhancing efficiency. Consequently, we compute the average total number of action expansions per layer across all problems, applying different methods within various domains, as a significant metric for comparison.
Total number of mutually exclusive actions. Mutually exclusive actions are generated from the set of candidate expansion actions based on many different mutually exclusive conditions. The critical aspect of this metric lies in the fact that a lower total number of mutually exclusive actions translates into significant computational time savings. This efficiency is evident both during the generation of the mutually exclusive action set in the forward expansion phase and in the filtering of candidate actions based on this set during the backtracking process.
Expansion Layers. Expansion layers hold significant importance in the GP algorithm, as each additional expansion layer exponentially increases the complexity of DFS during backtracking. Theoretically, the GP algorithm, provided it can compute a solution, achieves the minimal number of levels and the shortest action sequence. Therefore, maintaining a minimal increase in expansion levels is pivotal and it ensures the consistent attainment of a near-optimal action sequence.
Number of nodes for backtracking DFS. This metric serves as the cornerstone for validating our optimization efforts, as the DFS during backtracking accounts for the majority of the computational load in the GP algorithm, overshadowing the forward expansion phase. Particularly when dealing with increasing expansion depths, the exponentially growing number of DFS poses the most significant challenge for GP algorithms in tackling large-scale problems or complex solution sequences. We primarily utilize this metric to ascertain which method truly enhances the efficiency of the planning process.
Regarding the latter four metrics, our analysis was confined to data from the GP and LLMs4Plan-GPT4 methods, primarily for two reasons. Firstly, these metrics are relevant only in scenarios where the problem is successfully solved; failed solutions do not yield countable data. Consequently, we excluded LLMs4Plan-GPT3.5 from our statistical analysis due to its comparatively lower success rate. Secondly, these metrics are inherently calculable within the GP framework alone. Hence, directly solving problems using GPT-3.5 and GPT-4 precludes the possibility of gathering this data, as these methods operate outside the GP framework.
We display the above metrics in two distinct tables. Table 1 shows the problem-solving success rates, while table 2 focuses on the last four metrics previously mentioned.
Table 1: Success rate results. In the table, each row corresponds to a distinct domain, while each column represents a separate approach or method. The values presented within the table indicate the success rate for each combination of domain and approach, with these rates quantified on a scale ranging from 0 to 1.
GPT-3.5 GPT-4 GP LLMs4Plan-GPT3.5 LLMs4Plan-GPT4
gripper 0.0 0.6 0.6 0.0 0.6
miconic 0.1 0.5 0.3 0.1 0.5
logistics 0.2 0.6 0.6 0.2 0.8
movie 1.0 1.0 1.0 1.0 1.0
Table 2: LLM pruning results. In the table, each column signifies a distinct domain. It displays the statistical outcomes of the two methods across four indicators. Smaller numbers in the table denote lower computational resource usage and greater effectiveness.
Expansion Actions Mutex Actions Expansion Layers DFS Nodes
Domain GP LLMs4Plan GP LLMs4Plan GP LLMs4Plan GP LLMs4Plan
gripper 157 128 1656 1179 4 4 21321 498
miconic 181 176 792 668 5 6 2018484 11891
logistics 294 167 1726 161 5 5 1261250 59
movie 401 308 7 7 2 2 10869160 975163
5.3 Ablation Experiment
We conducted four ablation experiments to ascertain the effectiveness of forward pruning and backward sorting, detailed as follows:
  1. 1.
    LLMs4Plan-pruned-sorted: This method involves both forward pruning and backward sorting, corresponding to the LLMs4Plan-GPT4 approach in our experiments setup.
  2. 2.
    LLMs4Plan-pruned-unsorted: Here, we implement pruning without sorting.
  3. 3.
    LLMs4Plan-unpruned-sorted: In this approach, sorting is used, but not pruning.
  4. 4.
    LLMs4Plan-unpruned-unsorted: This method involves neither pruning nor sorting, akin to the GP method in our experimental setup.
Table 3 reveals that both pruning and sorting contribute to enhanced search efficiency, with their combination amplifying this effect. Notably, pruning appears slightly more effective than sorting. This is likely because LLM, while pruning, also organizes the remaining actions logically. In contrast, sorting may lead to minor errors due to the multitude of actions and lengthy text. The experiments also indicate that LLM tends not to prioritize empty actions in graph planning, favoring their later arrangement. This aligns with our analysis suggesting that prioritizing non-empty actions is more productive, as a layer without any action is essentially redundant.
Table 3: In the table, each row corresponds to a distinct domain, while each column represents a group of ablation experiments. The values presented within the table indicate the number of nodes for backtracking DFS. Both pruning and sorting effectively enhance search efficiency, leading to a substantial reduction in the number of nodes required for searching. Generally, pruning tends to be slightly more effective than sorting.
pruned-sorted pruned-unsorted unpruned-sorted unpruned-unsorted
gripper 498 1858 8317 21321
miconic 11891 47863 445145 2018484
logistics 59 85 1226 1261250
movie 975163 1211830 975163 10869160
5.4 Experimental Analysis
Analyzing Table 1, GPT-3.5 only handles simpler tasks like movie, showing low success rates in other domains and thus offering limited enhancement to the GP algorithm. GPT-4 exhibits significant improvements over GPT-3.5, particularly in reasoning and the decision of long action sequences. Leveraging GPT-4’s advanced reasoning and common sense, GP demonstrates improved problem-solving success in certain domains. The observed discrepancy largely stems from GP’s timeout failures in some cases. However, with LLM-based pruning, GP successfully resolves these issues within the set time limit, thereby boosting the success rate.
Turning to the second table, we assess how the GP method, augmented by LLM, enhances its success rate. The data indicates that both the total number of expansion actions and mutually exclusive actions experience a consistent decline with LLM integration. This suggests a reduction in the computational effort required for forward expansion. Concurrently, the number of expansion layers remains relatively stable, contributing to the generation of near-optimal action sequences. During the DFS in backtracking, LLM pruning significantly cuts down the number of search nodes, thereby endowing the GP method with expedited problem-solving capabilities, including for problems that were previously unsolvable.
We illustrate an example from the logistics domain in Figure 7, where we compare the number of expansion actions before and after LLM pruning. The application of LLM for pruning demonstrates significant efficacy across all layers.
Figure 7: An example of action pruning. Both the GP and LLMs4Plan-GPT4 methods expanded through 10 layers. The horizontal axis represents the layer number, with lower numbers indicating proximity to the initial state and higher numbers nearing the goal state. The vertical axis shows the count of expanded actions, including both domain-specific actions and numerous empty actions, characteristic of the graph planning algorithm. This implies a high pruning ratio for genuinely effective actions. LLM prunes almost every layer of expansion actions and the data in the table also contains many empty actions.
After checking the specific generated action sequences, we found that although the solution success rates of GPT-4 and our approach were almost equivalent, the length of the action sequences generated with GPT4 alone is generally longer. In contrast, integrating GPT-4 with graph planning yields more efficient and superior action sequences.
Upon analyzing examples where solutions failed, we observed that GPT4 is more prone to pruning errors at deeper expansion levels. This results in the discarding of effective actions, thereby unnecessarily increasing the expansion layers and hindering problem resolution. We attribute this to two primary factors. Firstly, as the expansion level deepens, both the predicate set and the candidate action set expand, leading to increasingly lengthy input prompts. This prolonged text can cause GPT-4 to gradually lose track of previous information, resulting in decision-making errors. Secondly, the nature of the predicate set in graph planning diverges from traditional planning’s current state representation. This discrepancy impairs LLM’s ability to accurately analyze the predicate set, leading to the erroneous elimination of effective actions. LLM lacks capacity to analyze complex predicate set combinations.
In response to GPT4’s pruning errors at deeper levels, we experimented with pruning only at shallower levels. This approach successfully avoids the earlier drawbacks, improving problem-solving success and efficiency. The key is that shallow-layer pruning with GPT-4 sidesteps the issues identified earlier. Therefore, identifying the optimal layers for LLM application in graph planning requires further research and experimentation.
Our analysis of additional failure examples indicates that graph planning excels in efficiently handling numerous non-mutually exclusive actions in parallel, due to its ability to group these actions within the same layer. However, its limitation becomes apparent in scenarios requiring the execution of longer, complex actions. If a problem’s optimal solution sequences are lengthy, the planning graph must be expanded considerably deeper. Despite effective pruning, this does not resolve the issue of exponential complexity growth in backtracking DFS caused by increased depth.
6 Conclusion and Future Work
Our comparative experiments in multiple domains demonstrated the efficacy of our LLMs4Plan in significantly enhancing the problem-solving capabilities of graph planning algorithms. Notably, LLMs4Plan boosts not just the success rate of problem resolution but also markedly enhances search efficiency and substantially reduces computational complexity. The runtime of LLMs4Plan is currently hindered by multiple LLMs calls. While our method requires multiple LLMs calls, it provides substantially improved results. There are also various ways to enhance runtime performance like using smaller LLMs like Llama (Touvron et al., 2023) or distilling LLMs’ knowledge into a smaller model (Shridhar et al., 2023; Hsieh et al., 2023; Liang et al., 2023). Those are interesting avenues for future research. Instead of leveraging LLMs to assist planning, it would also be possible to study acquring action models (Zhuo et al., 2011; Zhuo, 2015; Zhuo & Kambhampati, 2017) and more planning frameworks (Jin et al., 2022a) with the help of LLMs.